# Dynamic Order Template Prediction for Generative Aspect-Based Sentiment Analysis

### v1

@@ 4 Conclusion
We propose Dynamic Order Template (DOT) method for aspect-based sentiment analysis, addressing inefficiencies and out-of-distribution errors in static multi-view prompting. By dynamically generating necessary views based on instance-level entropy, DOT efficiently predicts the sentiment tuple in each instance. Our experiments on ASQP and ACOS datasets demonstrate that DOT achieves state-of-the-art F1-scores with reduced inference time, effectively balancing the strengths of single and multi-view approaches for ABSA.
@@
Also, since we directly connect first stage and second stage, if any errors occur, the errors may propagate and magnify as it moves to the subsequent stage. However, by splitting the task of ’predicting the appropriate number of tuples’ into two sub-tasks—’predicting the appropriate number of tuples’ and ’accurately predicting the tuples’—it becomes significantly easier to achieve accurate results in both areas, thereby enhancing overall performance in our work.
@@ Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
Appendix A Detailed Experimental Setups
We use AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 1e-4 for training two T5 models. We set the batch size to 16 for training and 24 for inference. We train the first stage model for 30 epochs, and train 40 epochs for the second stage. Additionally, we observe that the label of the datasets (i.e. sentiment tuples) irregularly contains stop words. For example, as in the first example of Figure 4, the inclusion of negations in the opinion terms is inconsistent. Also, as in the second example, element tuples sometimes contain ambiguous and meaningless stop words as an element. As a result, the fine-tuned model sometimes generates sentiment tuples containing stop words irregularly. It can yield critical performance degradation, even though they don’t affect the meaning of the sentiment elements. To resolve the problem from stop words, we filter these stop words using nltk package(Farkiya et al., 2015) for both generated results and dataset labels. We use four RTX 4090 GPUs to train and evaluate all of the models.
@@ Appendix A Detailed Experimental Setups
Appendix B Depth Analysis on First Stage
B.1 Accuracy on the Number of Views
We assess the accuracy of predicting the correct number of views in the first stage and present the results in Table 4. We evaluate the output by comparing it to the labeled sentiment tuples using RMSE and accuracy. We carefully implement the first stage baselines to compare our method properly as follows: Random: We find that the number of sentiment tuples in training dataset is mostly in range of 1 to 6. For each inference, we randomly sample one of the 6 numbers and compare it with our first stage result. Majority: We also reveal that about 60 percent of labels consist of single tuple. We construct a baseline that predicts only 1 for the number of tuples, to check whether our model has ability to predict the number of sentiment tuples of a sentence. Classification: We adopt the RoBERTa model Liu et al. (2019) to evaluate the results when treating the prediction of the number of views as a sequence classification task. We set the classes based on the number of sentiment tuples. As shown in 5, the distribution of tuple counts is skewed towards the lower end, with instances containing more than seven tuples being nearly non-existent. Consequently, we limit the categories from 1 to 6 and clip instances with 7 or more tuples to 6. Additionally, to address label imbalance, we employ a weighted loss function, where the weights are set as the inverse of the frequency ratio for each category as in Equation (4). We use same notation as in Section 2.1, and $\mathcal{I}\left(\right)$ denotes indicator function. This approach enables the model to effectively classify even the less represented classes.
@@ B.1 Accuracy on the Number of Views
1111222233334444555566667777000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8number of sentiment tuplesfrequency ratioASQP-Rest15ASQP-Rest16ACOS-Lap16ACOS-Rest16
@@
B.2 Effect of Element Exclusions
@@
As in Table 4, our proposed method outperforms the other baselines and nearly predicts the actual distribution of sentiment tuples within a small margin of error. This result justifies the direct use of the output from the first stage in the second stage. Also, the overall performance improved as the difficulty of predicting the excluded marker token’s element increased. The results from the first stage do not appear to be strongly correlated with those of the second stage. This emphasizes the importance of using a well-tuned initial model for the second stage.
Appendix C Computing Inference Time
@@
200200200200400400400400600600600600800800800800001,00010001{,}0001 , $0002{,}$  [^0] Dataset SizeInference Time (s)Single-viewDynamic-viewMulti-view
@@
Appendix D Detailed Setups for LLM Experiments
@@
As in Table 1, we perform the ABSA task using the GPT 3.5 Turbo, LLaMa-3 8B, and Mistral 7B models, and compared the results with our DOT model. For the GPT model, we utilize in-context learning Brown et al. (2020). We randomly sample 10 instances and combine them with instruction format, and add it as a prompt. For the other two open-source LLMs, we employ instruction tuning Wei et al. (2021) with the training dataset for fine-tuning, using the same instructions as in GPT prompts. To ensure stable model training during fine-tuning, we utilize the LoRa Hu et al. (2021). We present the specific prompts and framework in Figure 7.
@@
Figure 7: Instruction format for two LLM frameworks. We utilize in-context learning for ChatGPT inference, and instruction-tuning for LLaMa-3 and Mistral inference respectively.

### v2

@@ 4 Conclusion
We propose Dynamic Order Template (DOT) method for generative ABSA, addressing inefficiencies and out-of-distribution errors. Experiments on nine datasets demonstrate that DOT achieves state-of-the-art performance with reduced inference time, effectively balancing the strengths of previous single and multi-view approaches for ABSA.
@@
Also, since we directly connect first stage and second stage, if any errors occur, the errors may propagate and magnify as it moves to the subsequent stage. It results in relatively large standard deviation for different seeds as reported in Table 10. However, by splitting the task of ’predicting the appropriate number of tuples’ into two sub-tasks—’predicting the appropriate number of tuples’ and ’accurately predicting the tuples’—it becomes significantly easier to achieve accurate results in both areas, thereby enhancing overall performance in our work.
Finally, we define the number of necessary views as the number of sentiment tuples for simplicity and efficiency. A more complex yet refined method for determining the necessary number of views could be further explored in future research.
@@ Acknowledgement
This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
Appendix A Detailed Experimental Setup
We use AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 1e-4 for training two T5 models. We set the batch size to 16 for training and 24 for inference. We train the first stage model for 30 epochs, and train 40 epochs for the second stage.
@@ Appendix A Detailed Experimental Setup
We also note that the dataset labels (i.e., the sentiment tuples) sporadically include stop words. For example, as in the first example of Figure 4, the inclusion of negations in the opinion terms is inconsistent. Also, as in the second example, element tuples sometimes contain ambiguous and meaningless stop words as elements. As a result, the fine-tuned model sometimes generates sentiment tuples containing stop words irregularly. It can yield critical performance degradation, even though they don’t affect the meaning of the sentiment elements. To resolve the problem of stop words, we filter these stop words using nltk package(Farkiya et al., 2015) for both generated results and dataset labels. We use four RTX 4090 GPUs to train and evaluate all of the models.
Appendix B Case Study
Figure 5: Case study for the three main types of results. Blue one denotes correct, red one denotes incorrect, and the yellow one denotes irrelevant.
The second shows an example that predicts requires fewer views, but the example actually requires more views. Our analysis reveals that such cases frequently occurs with implicit $O$. As shown in Table 1, this suggests that our model’s performance might lag behind other baselines on the ACOS Rest16 dataset, which contains many samples with implicit $A$ and $O$. Additionally, the model struggles with predicting infrequent $C$ in the training set. Incorporating the concept of self-information and defining the necessary number of views based on the ’amount of information in a sample’ could effectively address this issue.
The final case involves cases with multiple sentiment tuples and longer lengths. We explain that errors in this scenario stem from two main reasons. Firstly, longer sentences include extended phrases that modify $A$ or $O$. Including all these modifiers as elements often leads to errors, a common problem across different models that requires an alternative solution. Secondly, errors occur when the number of tuples is incorrectly predicted in the first stage. If the predicted number of tuples is insufficient, some target sentiment tuples might be overlooked. Conversely, overestimation leads to the extraction of irrelevant aspects, as depicted in the Figure 5. However, we optimize the first stage to reduce tuple count errors, which helped mitigate performance drops by minimizing incorrectly generated or overlooked tuples.
Appendix C Depth Analysis on First Stage
C.1 Accuracy on the Number of Views
We assess the accuracy of predicting the value of $\hat{K}$ and present the results in Table 4. We evaluate the output by comparing it to the number of labeled sentiment tuples using RMSE and accuracy. We carefully implement the first stage baselines to compare our method properly as follows: Random: We find that the number of sentiment tuples in the training dataset is mostly in the range of 1 to 6. For each inference, we randomly sample one of the 6 numbers and compare it with our first stage result. Majority: We also reveal that about 60 percent of labels consist of a single tuple. We construct a baseline that predicts only 1 for the number of tuples, to check whether our model has the ability to predict the number of sentiment tuples of a sentence. Classification: We adopt the RoBERTa model Liu et al. (2019) to evaluate the results when treating the prediction of the number of views as a sequence classification task. We set the classes based on the number of sentiment tuples. As shown in Figure 6, the distribution of tuple counts is skewed towards the lower end, with instances containing more than seven tuples being nearly non-existent. Consequently, we limit the categories from 1 to 6 and clip instances with 7 or more tuples to 6. Additionally, to address label imbalance, we employ a weighted loss function, where the weights are set as the inverse of the frequency ratio for each category as in Equation (4). We use the same notation as in Section 2.1, and $\mathcal{I}\left(\right)$ denotes the indicator function. This approach enables the model to effectively classify even the less represented classes.
@@ C.1 Accuracy on the Number of Views
11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8number of sentiment tuplesfrequency ratioASQP-Rest15ASQP-Rest16ACOS-Lap16ACOS-Rest16
@@
C.2 Effect of Element Exclusions
@@
As in Table 4, our proposed method outperforms the other baselines and nearly predicts the actual distribution of sentiment tuples within a small margin of error. This result justifies the use of the output from the first stage in the second stage. The first stage results in Table 4 do not exhibit significant performance differences among various exclusion. However, for the second stage results, which serve as the final output of this task, we observe a significant performance difference. The performance in the second stage is generally higher when O is omitted because generating O correctly is the most difficult and crucial task in quadruple prediction (Chebolu et al., 2023). If O is not trained in the first stage and is reused in the second stage, the model appears to focus more on learning about O compared to other elements, which already have some level of information.
Appendix D Computing Inference Time
@@
20020020040040040060060060080080080001,0001{,}0001 , 0002,0002{,}0002 ,  [^0] Dataset SizeInference Time (s)Single-viewDynamic-viewMulti-view
@@
Appendix E Input and Target Examples for Each Stage
In Figure 8, we provide detailed examples for input and output pairs in each stage. The input sentences in the dataset are presented in a basic sentence structure, while the labels consist of lists of sentiment tuples. To preprocess this data, during the first stage, the original input sentence is kept unchanged, and the target is set as the initial order template, which consisted of a number of views corresponding to the number of sentiment tuples in the label. In the second stage, the input is processed by appending the final order template as a prompt to the original input sentence. The target is then constructed by adjusting the order of the elements within the sentiment tuples to align with the corresponding views in the order template.
Figure 8: Examples for input and target from original dataset for both first and second stage.
Appendix F Analysis on Implicit Term
In Table 1, DOT suffers from predicting sentiment tuples in Rest and Clothing domains. We noted that the ACOS dataset contains a significant number of instances with implicit aspects or implicit opinions. Additionally, we discovered that the Rest and Clothing dataset are smaller in scale compared to other ACOS datasets. The scale of each dataset and the number of instances containing implicit terms are recorded in Table 5. Based on these observations, we hypothesized that the size of the dataset and the distribution of implicit terms contribute to the performance degradation observed in the Rest and Clothing datasets.
Datasets ASQP ACOS MEMD
R15 R16 Lap Rest M-Rest M-Laptop Books Clothing Hotel
total samples 834 1264 2934 1530 3622 2863 2092 1674 2481
implicit samples 272 446 1826 822 1801 1751 1523 1083 1278
implicit sample % 32.6 35.3 62.2 53.7 49.7 61.2 72.8 64.7 51.5
Table 5: The size of each dataset and the number of samples containing implicit terms. For ease of comparison, We also provide the percentage of samples with implicit terms relative to the total number of samples. It is evident that the implicit term ratio in the ACOS dataset is higher compared to that in the ASQP dataset.
Methods Rest M-Rest ¼ M-Rest ½ M-Rest full
Paraphrase 50.06 40.26 47.77 49.09
DOT 44.84 35.74 47.81 49.49
Table 6: F1 scores only for samples containing implicit terms. We report the performance in Rest dataset and the performance trends across different dataset scales.
As shown in Table 6, it is evident that the F1 score for instances containing implicit terms in the Rest dataset is significantly lower compared to using the paraphrase method. Additionally, we observed a performance degradation when training on a randomly selected quarter of the M-Rest dataset. However, as the amount of training data from the M-Rest dataset increased, the performance on implicit terms improved, eventually surpassing the F1 score of the paraphrase method in the full M-Rest dataset. This result demonstrates that the small size of the dataset with a high proportion of implicit terms is the primary cause of the performance degradation in the Rest and Clothing dataset. It also suggests that the performance is likely to improve as the dataset size increases.
Appendix G Additional Analysis
In this section, we conduct an in-depth analysis of various aspects of our model. For a comprehensive evaluation, we use Paraphrase and MvP as baselines, running identical experiments for comparison. We assess performance across multiple tasks using several benchmarks, including R15, R16, Lap, Rest, and M-Rest.
Different Backbone Model
We conduct the experiment using different encoder-decoder based model, BART Lewis (2019) as backbone model. We utilize BART with the same hyperparameters and data processing techniques applied to the T5 model for three methods including ours. However, as in Table 7, we observe a noticeable decline in overall F1-scores for all models, primarily due to insufficient hyperparameter tuning compared to T5. Nevertheless, as shown in the results, our method still outperforms the baseline models with BART, suggesting that its effectiveness is not highly dependent on the choice of backbone model.
Methods ASQP ACOS
R15 R16 Lap Rest M-Rest
Paraphrase 31.77 38.15 30.98 36.65 35.74
MvP 33.48 41.01 32.57 40.40 40.30
DOT 35.98 41.73 33.12 39.61 40.91
Table 7: F1 score on benchmark datasets using BART as the backbone model.
Complex Sentences
As mentioned in Appendix B, processing long and complex contexts is a well-known challenge, and our model performs similarly to others in this regard. We define complex sentences as those containing more than three sentiment tuples, exceeding 22 words in length, or having a Flesch-Kincaid Grade Level Solnyshkina et al. (2017) of 9 or higher, representing the top 20% for each criterion. We sample these complex sentences and evaluate the F1 scores for these samples. As shown in the Table 8, performance degradation in complex sentences is a common issue across all models. We attribute the larger performance drop in our model compared to MvP to the fact that it uses fewer views, which limits its capacity to thoroughly analyze complex sentences. Nevertheless, our model’s final Complex F1 score remains close to that of MvP and surpasses that of the Paraphrase model.
Methods ASQP ACOS
R15 R16 Lap Rest M-Rest
Paraphrase 44.94 55.73 37.15 56.12 54.37
MvP 46.71 58.00 37.68 56.06 58.23
DOT 46.93 57.70 38.57 54.16 57.72
Table 8: F1 scores evaluated on complex samples only.
Training Complexity
Our method may appear complex due to the numerous components that require training. However, since our method involves simply training the T5 model twice without complex optimization procedures, the overall training time is not significantly longer than that of other models. As shown in the Table  9, even though our model uses 30 and 40 epochs for two stages of training—more than the 20 epochs used in MvP—the total training time remains much shorter than that of MvP. In terms of memory usage, only two T5 models are allocated in memory, so the memory consumption does not increase exponentially compared to existing models.
Methods ASQP ACOS
R15 R16 Lap Rest M-Rest
Paraphrase 212.83 314.18 652.31 349.79 815.48
MvP 3883.74 5008.84 11006.07 6169.02 14634.71
DOT 1161.73 1648.61 3310.63 1814.41 4157.93
Table 9: Training duration for each benchmark.
Standard Deviation
We conduct experiments using five different random seeds and calculate the standard deviation of the outcomes. Results are reported in Table 10. Our findings indicate that our model exhibits a higher overall standard deviation compared to other baselines. This can be attributed to the structure of the method, where an error at one stage is likely to propagate and accumulate. However, it is important to note that the absolute value of the standard deviation is not significantly large. In fact, the higher variation suggests that the model may possess greater potential to achieve stronger performance.
Methods ASQP ACOS
R15 R16 Lap Rest M-Rest
Paraphrase $\pm$ 0.44 $\pm$ 0.64 $\pm$ 0.26 $\pm$ 0.68 $\pm$ 0.38
MvP $\pm$ 0.54 $\pm$ 0.29 $\pm$ 0.48 $\pm$ 0.72 $\pm$ 0.48
DOT $\pm$ 0.74 $\pm$ 0.85 $\pm$ 1.01 $\pm$ 0.76 $\pm$ 0.42
Table 10: Standard deviation of outcomes in Table 1.
Appendix H Detailed Setups for LLM Experiments
@@
As in Table 1, we perform the ABSA task using the GPT-4o, LLaMa-3.1-8B, and Mistral-7B models, compairing the results with our DOT model. For the GPT model, we utilize in-context learning Brown et al. (2020). We randomly sample 10 instances and combine them with instruction format, and add it as a prompt. For the other three open-source LLMs, we employ instruction tuning Wei et al. (2021) with the training dataset for fine-tuning, using the same instructions as in GPT prompts. To ensure stable model training during fine-tuning, we utilize the LoRa Hu et al. (2021). We present the specific prompts and framework in Figure 9.
@@
Figure 9: Instruction format for two LLM frameworks. We utilize in-context learning for GPT-3.5-turbo inference, and instruction-tuning for LLaMa-3.1 and Mistral inference respectively.